What is the wavelength of electromagnetic radiation that has a frequency of 6.1 x 10(To the 18th power) S(with the power of -1)? Show steps.


1
Expert's answer
2020-05-08T14:09:59-0400

Wavelength, λ = c/ frequency

Where c = speed of light

Here frequency =6.1×10( to the power 18) s(to the power - 1)

c = 3× 10(to the power 8) meter /s

Wavelength, λ=

(3×10(to the power 8))/ (6.1×10(to the power 18)) meter

=0.5 × 10(to the power-10) meter